OpenStudy (anonymous):

Numbers 1- 10 are written on cards and placed in a bag. Find the probability of choosing an 8 or choosing a number less than 5. Enter your answer in simplified fraction form; example: 3/5.

OpenStudy (anonymous):

Choosing an 8: 1/10. Choosing a number less than 5: 2/5

OpenStudy (anonymous):

Thanks :) @justin12lo

OpenStudy (kropot72):

The events 'choose an 8' and 'choose a number less than 5' are mutually exclusive. Therefore the required probability is the sum of the individual probabilities: \[P(8\ or\ <5)=\frac{1}{10}+\frac{4}{10}=\frac{5}{10}=\frac{1}{2}\]